OpenStudy (anonymous):

simplify the square root of 24 by factoring

OpenStudy (jonnyvonny):

So, when anything is being square-rooted, you may take break apart a perfect square (4,9,16 ect), within the radical, reduce that, then multiply with the remainder. Like so:\[\sqrt{24}\rightarrow \sqrt{4}*\sqrt{6}\rightarrow2*\sqrt{6}\rightarrow2\sqrt{6}\] I hope this helped, if you need further elaboration, another example, or help with anything else, feel free to pst. :)

OpenStudy (anonymous):

it did thank you